Could you please assist me with this problem? The directions are: solve each equation. State the number of roots and the type of roots.
4X^2 + 1=0

Answer by TimothyLamb(4379)

---
4x^2 + 1 = 0
---
the above quadratic equation is in standard form, with a=4, b=0, and c=1
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
4 0 1
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two complex roots at:
---
x = 0 + (0.5)i
x = 0 - (0.5)i
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php